Find the curvature and torsion of the curve x=a(u-sinu),y=(1-cosu),z=bu


1
Expert's answer
2021-06-07T18:44:47-0400

let

"r=[a(u-sinu),1-cosu,bu]"

so

"r^{'}" (single deriverative of r) "=[a(1-cosu),sinu,0]"

"r^{''}=[asinu,cosu,0]"

"r^{'''}=acosu,-sinu,0"


Now :


Curvature "=\dfrac{|r^{'} \times r^{''}|}{|r^{'}|^3}"


Now: "r^{'} \times r^{''}=" "i[-bcosu]-j[-absinu]+k[a(cosu-cos^2u)-asin^2u]"

"=(-bcosu)i +(absinu)j+(acosu-a)k"


"|r^{'} \times r^{''}|= \sqrt{b^2(cos^2u+a^2sin^2u)+a^2+a^2(cos^2u-2cosu)}"


"|r^{'}|= \sqrt{a^2(1+cos^2u)-2acosu+sin^2u+b}"


Curvature "= \dfrac{|r^{'} \times r^{''}|}{|r^{'}|^3} = \dfrac{\sqrt{b^2(cos^2u+a^2sin^2u)+a^2+a^2(cos^2u-2cosu)}}{(a^2(1+cos^2u)-2acosu+sin^2u+b)^{3/2}}"


Torsion "= \dfrac{[r^{'},r^{''},r^{'''}]}{|r^{'} \times r^{''}|^2}"


"[r^{'},r^{''},r^{'''}]=[a(1-cosu)(0)] -[sinu(0)] +b(-asin^2u-acos^2u)= -ab"


so Torsion "= \dfrac{-ab}{b^2(cos^2u+a^2sin^2u)+a^2+a^2(cos^2u-2cosu)}"